OpenStudy (anonymous):

what aret the zeros of the function f(x) = x^2 - 5x

OpenStudy (anonymous):

The zeros of a function are when the function is equal to zero. If you graph that function, the zeros will be whenever it crosses the x-axis. But first, let's set f(x) equal to 0: f(x) = 0 = x^2 - 5x. We can factor out an 'x': 0 = x(x-5). This says that either 0 = x or (x-5) = 0. If we solve for 'x' in the second equation, we'll have x=5. So, the zeros must be x=0 and x=5.